OpenStudy (anonymous):

–5x + 3y = 6 x – y = 4

myininaya (myininaya):

have you tried to solve this by elimination

myininaya (myininaya):

I would multily the bottom equation by 3 and add the equations together after that this will result in the elimination of y and you will be able to solve for x first

OpenStudy (anonymous):

so is (–9, –13)

OpenStudy (triciaal):

following the directions above to solve by elimination, (-5+3 ) x =6 + 12 x = 18/-2 = -9 y = x + 4 = -9 + 4 = -5

myininaya (myininaya):

y=x-4=-9-4=-13

OpenStudy (triciaal):

check in –5x + 3y = 6 -5(-9) + 3 (-5) = 45-15 = 30 error @myininaya thanks I see my error

myininaya (myininaya):

We all make them. I make them more than most.

OpenStudy (triciaal):

LHS: -5(-9) + 3(-13) = 45 -39 =6 = RHS